# Robust Plan Evaluation based on Approximate Probabilistic Machine Learning

### v2

@@ Abstract.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
@@ 1. Introduction
Executing queries in database management systems involves accessing, joining, and aggregating data from various sources. These queries have multiple execution plans, with significant variations in performance. To find the best plan at compile-time, sophisticated query optimizers use limited search algorithms and a cost model. The cost model evaluates each plan based on parameters like data size and cardinalities, which are estimated since they are unknown at compile-time. Moreover, the cost model makes certain assumptions that may not hold in practice. Consequently, the plan selected by the optimizer may be suboptimal at runtime, lacking robustness against estimation errors and non-conforming environments. Query optimization approaches that are less sensitive to estimation errors and do not rely on simplifying assumptions are considered robust. This work focuses on the robustness in the context of plan optimization. Robustness in other phases of query optimization and processing, like query rewrite and query execution, is beyond the scope of this research. While various approaches are proposed in the literature for robust query optimization, they are rarely adopted in practice due to their limitations. Some require intrusive changes in the query execution engine (Babu et al., 2005; Markl et al., 2004; Moumen et al., 2016), have unreasonable compilation overhead (Dutt and Haritsa, 2016; Karthik et al., 2016, 2018), or make unrealistic assumptions (Chu et al., 2002; Wolf et al., 2018). With the advent of new approaches based on ML, some of these limitations are mitigated. For example, Neo (Marcus et al., 2019) shows some robustness to misestimation of the input parameters, and Bao (Marcus et al., 2024) demonstrates minimal regressions for the tested workloads. However, these approaches too either do not quantify uncertainties (Chen et al., 2023; Marcus et al., 2024, 2019) or do not use uncertainties to effectively improve a robustness measure (Yu et al., 2022). In addition, ML-based approaches suffer from significant errors when there’s a drift in the data or workload distributions. While ML-based techniques show promising advantages for reducing the maintenance and tuning overheads of query optimizers, a successful transition to such approaches relies on addressing concerns about their robustness and proposing methods for minimizing the risk of regressions. Given the limitations in prior work, a practical approach to solve the robustness problem needs to a) have limited compilation overheads, b) avoid making simplifying assumptions, c) formalize and quantify the notion of robustness, d) quantify uncertainties associated with cost estimates, and e) use uncertainty estimates effectively to improve a robustness measure. We propose Roq, a novel approach for robust query optimization using approximate probabilistic ML. Our contributions are:
  • Formalizing the problem of robust query optimization in two different ways
@@
  • Formalizing the notion of plan and cost model robustness to enable an objective evaluation of these characteristics in this work as well as in future research.
@@
  • Proposing theoretical methods for quantifying uncertainties and risks in the context of plan evaluation and selection based on approximate probabilistic ML.
@@
  • Proposing a new model architecture for a learned cost model that supports measuring uncertainties and provides improved performance compared with the state-of-the-art.
@@
  • Proposing novel strategies and algorithms for plan evaluation and selection that outperform the state-of-the-art approaches by accounting for uncertainties and risks.
@@
  • Demonstrating the robustness of the proposed approaches to workload shifts and the limited compilation overheads which makes them suitable for practical use.
@@
  • An ablation study of the impacts of accounting for different types of risk in the proposed risk-aware strategies.
@@ 2. Problem Statement
Query optimization involves exploring a large space of candidate execution plans, recursively evaluating a set of plan fragments (sub-plans), selecting the one with the minimum expected cost, and constructing larger plans from smaller fragments, until the creation of complete plan. In comparing plan fragments query optimizers do not consider the uncertainty of cost estimations, as they assume that the costs, as estimated by the cost model, have zero variance. As a result, query optimizers target expected optimality, but not robustness. In this paper, when we attribute a characteristic to current query optimizers, we refer to Db2 to represent the state-of-the-art commercial DBMSs. Although we recognize the differences among the optimization approaches of data systems, to the best of our knowledge, the limitations impacting their robustness are prevalent in the major ones. In the following, we discuss intuitively the robustness problem and we present a formal definition.
@@ 2.1. Problem Discussion
Uncertainty of sub-plan evaluation and comparison is traced back to 3 main sources: uncertainty rooted in the (a) plan structure and the data that flow through the plan nodes, (b) limitations of the estimation method, and (c) a crude comparison of a set of plans. We use the terms risk and risky to point out a high level of uncertainty in each of these dimensions. We use the term robust if the sensitivity to uncertainty is low. Hence, a plan can be considered robust if its execution cost is not sensitive to changes in the parameter estimates and the environment variables. The level of plan robustness is influenced by its structural characteristics, the operators used in the plan, the patterns of error propagation, and the characteristics of the underlying data. Similarly, an estimation method can be considered robust if it is not sensitive to inaccurate parameter estimates and non-conforming environments. Also, an estimation method should quantify the expected uncertainty for each estimate in order to support robustness. Classic cost models are not robust since they are sensitive to inaccurate parameter estimates and non-conforming environments and do not quantify estimation uncertainty. Due to the uncertainties rooted in the plan structure and data, and the limitations of the cost estimation method, picking a plan from a set of plans is also an uncertain task, as there is a risk that the selected plan will be suboptimal at runtime. We define the sources of uncertainty (or else, risk) in the plan evaluation and selection process as follows:
@@ 2.2. Problem Formulation
We first formulate the classical problem of optimal plan selection; based on this, we provide two alternative formulations of the problem of robust plan selection. The latter are employed to propose risk-aware plan selection strategies in Section 3.3.
Optimal plan selection problem: Given a finite set of candidate execution plans $\mathcal{P}=({p_{1},p_{2},\ldots,p_{n}})$, and a cost function $(f(p_{i}))$ that estimates the cost of executing a plan $(p_{i}),i=1,\ldots,n$, find the plan $(p^{*})$ such that: $p^{*}=\arg\min_{p_{i}\in\mathcal{P}}f(p_{i}).\square$
@@
Robust plan selection problem (Approach 1): Given a finite set of candidate execution plans $\mathcal{P}=({p_{1},p_{2},\ldots,p_{n}})$, a cost function $(f(p_{i}))$ and a robustness function $(g(p_{i}))$ that estimate the cost and robustness of a plan $(p_{i}),i=1,\ldots,n$ respectively, find the plan $(p^{*})$ such that: $p^{*}=\arg\min_{p_{i}\in\mathcal{P}}\left(SOR(p_{i})\right)$ where $SOR(p_{i})$ is the risk of suboptimality of the plan $p_{i}$ compared to the alternatives, defined as a function $k(.,.)$ of $f(p_{i})$ and $g(p_{i})$, i.e. $SOR(p_{i})=k(f(p_{i}),g(p_{i}))$. $\square$
[…]
@@
Robust plan selection problem (Approach 2): Given a finite set of candidate execution plans $\mathcal{P}=({p_{1},p_{2},\ldots,p_{n}})$, a cost function $(f(p_{i}))$ and a robustness function $(g(p_{i}))$ that estimate the cost and robustness of a plan $(p_{i}),i=1,\ldots,n$ respectively, find the plan $(p^{*})$ such that: $p^{*}=\arg\min_{p_{i}\in\mathcal{P}}\left(\alpha\cdot f(p_{i})+(1-\alpha)\cdot g(p_{i})\right)\square$
Solving the problem based on this formalization necessitates providing an approximation for the cost function $(\hat{f}(.))$, the robustness function $(\hat{g}(.))$, and a principled approach for determining a suitable weight parameter $\alpha$.
Most of the prior work has been concerned with providing a better approximation for the cost function $(\hat{f}(.))$ (Marcus et al., 2019, 2024; Chen et al., 2023; Zhang et al., 2023; Hilprecht and Binnig, 2022; Siddiqui et al., 2020). However, providing clear definitions and approximation methods for $\hat{g}(.)$ and $\hat{k}(.,.)$ are missing in the literature. In this paper, we aim at filling in this gap.
@@ 3. Theoretical Framework
In this section, we study the sources of uncertainty in the context of plan evaluation and selection towards formalizing the notion of robustness. Section 3.1 presents a study and definitions for the notions of plan robustness and cost model robustness, and forms guidelines for the design of a robust cost model. Section 3.2 presents quantification measures for the 3 types of risks tailored for an ML-based cost model. Section 3.3 presents plan evaluation and selection strategies and algorithms that employ risk measures to ensure robustness.
@@ 3.1. Studying and Modeling Robustness
The uncertainty of a plan cost estimate is rooted in two main sources: a) the plan structure and the data that flow through the plan operators and b) the limitations of the cost model used to evaluate the plan.
@@ 3.1.1. Plan and Cost Estimation Uncertainty Modeling
The robustness of a plan can be formalized based on the behavior of its cost given various parameter values(Wolf et al., 2018). The cost of a plan is usually represented by a Parametric Cost Function (PCF) that shows its cost at different points in a multidimensional parameter space, which includes all combinations of values of different error-prone parameters that impact the cost of a plan. Previous works have suggested various properties of the PCF to measure plan robustness. For example, the slope of the PCF and its area under the curve are proposed as such measures (Wolf et al., 2018).
We formalize this notion by modeling the uncertainty in the PCF and how it can be decomposed into uncertainties from various sources, including the structure of the plan and the parameters of the cost estimation model.
@@
(1) $\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}=\mathcal{X}^{*})=E[\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\theta)]+\text{Var}(E[f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\theta])$
where $\mathcal{X}^{*}$ represents the estimated distribution of the input parameters.
@@
The decomposition provided in Theorem 3.1 is illustrated in Figure 1, with a simple linear cost model with one error-prone cardinality parameter. The first term ($E[\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\theta)]$) has a direct relationship with a) the error in the input cardinality and b) the slope of $f_{\theta}(\mathcal{X})$ which determines its sensitivity to the error in the input. Therefore, its quantification is a suitable measure for plan robustness. The second term ($\text{Var}(E[f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\theta])$), on the other hand, is directly influenced by the uncertainty in the model parameters, and it is a suitable measure for model robustness. This represents the uncertainty rooted in our lack of knowledge about an ideal cost model. A proof for Theorem 3.1 is provided in Appendix A.1.
@@ 3.2. Ensuring Robustness via Risk Quantification
We recognized the probabilistic nature of cost estimates and provided guidelines for the creation of a robust cost model based on ML. We continue with the definition of quantification measures of the 3 types of risks, which will be employed to ensure robustness. The proposed quantification is tailored to an ML-based cost model and employs approximate probabilistic ML. We describe an example with scenarios of plan cost uncertainties that is used in the discussion of the proposed measures.
@@
Figure 2 illustrates scenarios for plan cost inaccuracies. Each plot shows the probability density function (PDF) of the estimated cost of two alternative plans. The costs are assumed to have a normal distribution. In each case, assuming the optimizer estimates the most likely cost for each plan, the plan with minimum expected cost is considered as the optimal one. The alternative plan, while having a higher expected cost, is also likely to be optimal at runtime with a non-zero probability. Note that the best-case scenario for the optimizer is when the plan costs are consistently accurate, as the optimizer naively assumes (Figure 2a). In such a case, the likelihood of plan X being more expensive than plan Y is low. More likely, however, the estimates have various levels of inaccuracies. Figure 2b/c/d represent such scenarios. If one of the two estimates is relatively more inaccurate, the risk can increase substantially (Figure 2b/c). The worst-case is if both estimates are relatively inaccurate (Figure 2d), as the risk of choosing the expected optimal plan increases. This example highlights the importance of taking into account the PDFs of the cost estimates rather than only the expected values. $\square$
@@ 3.2.1. Uncertainty Modeling in ML
(3) $\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}=\mathcal{X}^{*})=E[\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\mathcal{D})]+\text{Var}(E[f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\mathcal{D}])$
The first term in Equation 3 is aleatoric (or data) uncertainty and the second term is epistemic (or model) uncertainty (Kendall and Gal, 2017). The variability of the behavior of the real-world phenomena is explained by data uncertainty. Data uncertainty can also be caused by noisy input vectors or labels, as well as low-dimensional input vectors that do not adequately explain the sample (Hüllermeier and Waegeman, 2021). Model uncertainty, however, captures the variability of predictions due to a lack of knowledge about an ideal model. This is rooted in limitations in model architecture, training procedure, or the coverage of the training data (Gawlikowski, 2023). For a learned cost model, these notions can be quantified using various techniques from the ML literature (Gawlikowski, 2023). In the following sections, we explain the approaches we take to quantify these two types of uncertainty.
Data Uncertainty A neural network can be designed to predict the parameters of probability distribution (Nix and Weigend, 1994). i.e. it not only predicts the conditional expected value but also the conditional variance of the target given the training data and the input sample ($E[Var(y|x^{*},D)]$). This is done by adding a second branch to the output of the neural network, which predicts the variance. Optimizing the following loss function corresponds to maximizing the log-likelihood with a Gaussian prior as shown in (Nix and Weigend, 1994):
(4) $\text{Loss}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\log(\sigma_{i}^{2})}{2}+\frac{(y_{i}-(\mu_{i})^{2})}{2(\sigma_{i}^{2})}+\frac{1}{2}\log 2\pi\right)\quad$
where $(\mu_{i})^{2}$ is plugged in from the first branch of the output layer, $(\sigma_{i}^{2})$ from the second one, and $y_{i}$ from the labels.
Model Uncertainty Model uncertainty is captured by the variance of the conditional expected values of the target ($Var(E[(y|x^{*},D)])$). To capture this, we use approximate variational inference by Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016). In this method, the neural network is trained by applying dropout on hidden layers with a Bernoulli distribution. At inference time, dropout is again applied randomly to the hidden layers, resulting in variations in the weight matrices, hence in the predicted values. Note that with a Gaussian prior, this corresponds to combining multiple Gaussian distributions $\{N(\mu_{\hat{\theta}_{t}},\sigma_{\hat{\theta}_{t}}^{2})\}_{t=1}^{T}$ into one $N(\hat{\mu}_{t},\hat{\sigma}_{t}^{2})$. Given each draw of model weights results in two estimated parameters, we have a set of parameter estimates $\{(\hat{\mu}_{t},\hat{\sigma}_{t}^{2})\}_{t=1}^{T}$. The predicted mean is:
@@
$\hat{\sigma}_{m}^{2}=\text{Var}\left(E[(y|x^{*},D)]\right)=E\left[(y|x^{*},D)^{2}\right]-\left(E\left[(y|x^{*},D)\right]\right)^{2}\$
(7) $=\frac{1}{T}\sum_{t=1}^{T}(\hat{\mu}_{t}^{2})-\left(\frac{1}{T}\sum_{t=1}^{T}(\hat{\mu}_{t})\right)^{2}$

### v3

@@ Abstract.
Proceedings of the VLDB Endowment, Vol. 18, No. 8 ISSN 2150-8097.
@@ 1. Introduction
Executing queries in database management systems involves accessing, joining, and aggregating data from various sources. These queries have multiple execution plans, with significant variations in performance. To find the best plan at compile-time, query optimizers use limited search algorithms and a cost model. The cost model evaluates each plan based on parameters such as cardinalities, which are estimated as they are unknown at compile-time. Moreover, the cost model makes certain assumptions that may not hold in practice. Consequently, the plan selected by the optimizer may be suboptimal at runtime, lacking robustness against estimation errors and non-conforming environments. Query optimization approaches that are less sensitive to estimation errors and do not rely on consequential simplifying assumptions are considered robust. This work focuses on the robustness in the context of plan optimization. Robustness in other phases of query optimization and processing, like query rewrite and execution, is beyond the scope of this research. While various approaches are proposed in the literature for robust query optimization, they are rarely adopted in practice due to their limitations. Some require intrusive changes in the query execution engine (Babu et al., 2005; Markl et al., 2004; Moumen et al., 2016), have unreasonable compilation overhead (Dutt and Haritsa, 2016; Karthik et al., 2016, 2018), or make unrealistic consequential assumptions (Christodoulakis, 1984; Chu et al., 2002; Wolf et al., 2018). With the advent of new approaches based on ML, some of these limitations are mitigated. For example, Neo (Marcus et al., 2019) shows some robustness to misestimation of the input parameters, and Bao (Marcus et al., 2024) demonstrates minimal regressions for the tested workloads. However, ML-based approaches suffer from significant errors when there is a drift in the data or workload distributions (Hilprecht et al., 2020). While ML-based techniques show promising advantages for reducing the maintenance and tuning overheads of query optimizers, a successful transition to such approaches relies on addressing concerns about their robustness. Given the limitations in prior work, a practical approach to solve the robustness problem needs to a) have limited compilation overheads, b) avoid making consequential or unjustified simplifying assumptions, and c) demonstrate robustness based on well-defined robustness measures. We propose Roq, a novel approach for robust query optimization using approximate probabilistic ML. Our contributions are:
  • We formalize the problem of robust query optimization.
@@
  • We formalize the notion of plan and cost model robustness to enable an objective evaluation of these characteristics.
@@
  • We formalize theoretical methods for quantifying uncertainties and risks in the context of plan evaluation and selection based on approximate probabilistic ML.
@@
  • We propose a new model architecture for a learned cost model that supports measuring uncertainties and provides improved performance compared with the state-of-the-art.
@@
  • We propose novel strategies and algorithms for plan evaluation and selection that outperform the state-of-the-art approaches by accounting for uncertainties and risks.
@@
  • We demonstrate the robustness of the proposed approaches to workload shifts and their limited compilation overhead.
@@
  • We conduct an ablation study of the impacts of accounting for different types of risk in the risk-aware strategies.
The proposed risk quantification and measurement techniques, along with the risk-aware strategies, are not tightly coupled with the proposed model architecture. Any NN-based learned cost model can benefit from employing these techniques to enhance performance and robustness with minimal overhead.
@@ 2. Problem Statement
Query optimization involves exploring a large space of candidate execution plans, recursively evaluating a set of plan fragments (sub-plans), selecting the one with the minimum expected cost, and constructing larger plans from smaller fragments, until the creation of a complete plan. In comparing plan fragments, query optimizers do not consider the uncertainty of cost estimations; i.e., they practically assume that the estimated costs have zero variance. Therefore, they target expected optimality, but not robustness. [^1] In this paper, when we attribute characteristics to query optimizers, we refer to that of Db2 as a representative of the state-of-the-art commercial RDBMSs. Although we recognize the differences among the optimization approaches of data systems, to the best of our knowledge, the limitations impacting their robustness are prevalent in the major ones. In the following, Section 2.1 discusses intuitively the robustness problem and defines types of risk. Section 2.2 provides formal problem definitions using two alternative approaches which are later used in the proposed plan selection strategies in Section 3.3.
@@ 2.1. Problem Discussion
Uncertainty in sub-plan evaluation and selection is traced back to three main sources: uncertainty rooted in the (a) plan structure and the data that flow through the plan nodes, (b) limitations of the estimation method, and (c) a crude comparison of a set of plans. We use the terms risk and risky to point out a high level of uncertainty in each of these dimensions. We use the term robust if the sensitivity to uncertainty is low. Hence, a plan can be considered robust if its execution cost is not sensitive to changes in the parameter estimates and the environment variables. The level of plan robustness is influenced by its structural characteristics, the operators used in the plan, the patterns of error propagation, and the characteristics of the underlying data. Similarly, an estimation method can be considered robust if it is not sensitive to inaccurate parameter estimates and non-conforming environments. Also, an estimation method should quantify the expected uncertainty for each estimate in order to support robustness. Classic cost models are not robust since they are sensitive to inaccurate parameter estimates and non-conforming environments, and do not quantify estimation uncertainties. Due to the uncertainties rooted in the plan structure and data, and the limitations of the cost estimation method, picking a plan from a set of plans is also an uncertain task, as there is a risk that the selected plan will be suboptimal at runtime. We define the sources of uncertainty (or else, risk) in plan evaluation and selection:
@@ 2.2. Problem Formulation
We first formulate the classical problem of optimal plan selection; based on this, we provide two alternative formulations of the problem of robust plan selection. The latter are the basis for the proposed risk-aware plan selection strategies presented in Section 3.3.
[…]
@@
Robust plan selection problem (Approach 1): Given a finite set of candidate execution plans $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$, a cost function $f(.)$ and a robustness function $g(.)$ that estimate the cost and robustness of any plan $\{p_{i},\forall i\in\{1,\ldots,n\}\}$ respectively, find the plan $p^{*}$ such that: $p^{*}=\arg\min_{p_{i}\in\mathcal{P}}SOR(p_{i})$ where $SOR(p_{i})$ is the suboptimality risk of the plan $p_{i}$ compared to the alternatives, defined as a function $k(.,.)$ of $f(.)$ and $g(.)$, i.e. $SOR(p_{i})=k(f(p_{i}),g(p_{i}))$. $\square$
This approach formalizes the problem as a risk minimization problem, where solving it requires providing approximations for the cost function $\hat{f}(.)$, the robustness function $\hat{g}(.)$ and a formulation for the suboptimality risk function $k(.,.)$. We use this formalization to propose a risk-aware plan selection strategy based on suboptimality risk in Section 3.3.1
@@
Robust plan selection problem (Approach 2): Given a finite set of candidate execution plans $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$, a cost function $f(.)$ and a robustness function $g(.)$ that estimate the cost and robustness of any plan $\{p_{i},\forall i\in\{1,\ldots,n\}\}$ respectively, find the plan $p^{*}$ such that: $p^{*}=\arg\min_{p_{i}\in\mathcal{P}}\left(\alpha\cdot f(p_{i})+(1-\alpha)\cdot g(p_{i})\right)\square$
Solving the problem based on this formalization necessitates providing an approximation for the cost function $\hat{f}(.)$, the robustness function $\hat{g}(.)$, and a principled approach for determining a suitable weight parameter $\alpha$.
Most of the prior work has been concerned with providing a better approximation for the cost function $\hat{f}(.)$ (Marcus et al., 2019, 2024; Chen et al., 2023; Zhang et al., 2023; Hilprecht and Binnig, 2022; Siddiqui et al., 2020). However, providing clear definitions and approximation methods for $\hat{g}(.)$ and $\hat{k}(.,.)$ are missing in the literature. In this paper, we aim at filling in this gap.
@@ 3. Theoretical Framework
We study the sources of uncertainty in the context of plan evaluation and selection towards formalizing the notion of robustness. Section 3.1 presents a study and definitions for the notions of plan robustness and cost model robustness, and forms a foundation for the design of a robust cost model. Section 3.2 presents quantification measures for the three types of risks tailored for an ML-based cost model. Section 3.3 presents plan evaluation and selection strategies and algorithms that employ risk measures to ensure robustness.
@@ 3.1. Studying and Modeling Robustness
The uncertainty of a plan cost estimate is rooted in two main sources: a) the plan structure and the data that flow through the plan operators and b) the limitations of the cost model.
@@ 3.1.1. Plan and Cost Estimation Uncertainty Modeling
Plan robustness can be formalized based on cost behavior given various parameter values (Wolf et al., 2018). The plan cost is represented by a Parametric Cost Function (PCF) in a multidimensional parameter space, which includes all combinations of values of different error-prone parameters that impact the plan cost. Previous works have suggested various properties of the PCF to measure plan robustness, such as the slope of the PCF and its area under the curve (Wolf et al., 2018). We formalize this notion by modeling the uncertainty in the PCF and how it can be decomposed into uncertainties from various sources, including the plan structure and the parameters of the cost model.
@@
(1) $\begin{split}\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}=\mathcal{X}^{*})&=E[\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\theta)]\\ &+\text{Var}(E[f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\theta])\end{split}$
where $\mathcal{X}^{*}$ is the estimated distribution of the input parameters.
@@
The decomposition provided in Theorem 3.1 is illustrated in Figure 1, with a simple linear cost model with one error-prone cardinality parameter. The first term ($E[\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\theta)]$) has a direct relationship with a) the error in the input cardinality and b) the slope of $f_{\theta}(\mathcal{X})$ which determines its sensitivity to the error in the input. Therefore, its quantification is a suitable measure for plan robustness. The second term ($\text{Var}(E[f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\theta])$), on the other hand, is directly influenced by the uncertainty in model parameters, and is a suitable measure for model robustness. This represents the uncertainty rooted in our lack of knowledge about an ideal cost model. A proof for Theorem 3.1 is provided in Appendix A.1.
@@ 3.2. Ensuring Robustness via Risk Quantification
We recognized the probabilistic nature of cost estimates and provided a decomposition model for their uncertainties. We continue with the definition of quantification measures of the three types of risks, which will be employed to ensure robustness. The proposed quantification is tailored to an ML-based cost model and employs approximate probabilistic ML. We describe an example with scenarios of plan cost uncertainties that is used in the discussion of the proposed measures.
@@
Figure 2 illustrates various scenarios of cost uncertainties. Each plot shows the probability density function (PDF) of the estimated cost of two alternative plans. The costs are assumed to have a normal distribution. In each case, assuming the optimizer estimates the most likely cost for each plan, the plan with the minimum expected cost is selected as the optimal. The alternative plan, while having a higher expected cost, is also likely to be optimal at runtime with a non-zero probability. Note that the best-case scenario for the optimizer is when the plan costs are consistently accurate, as the optimizer naively assumes (Figure 2a). In such a case, the likelihood of plan X being more expensive than plan Y is low. More likely, however, the estimates have various levels of uncertainty. Figure 2b/c/d represent such scenarios. If one of the two estimates is relatively more uncertain, the risk can increase substantially (Figure 2b/c). The worst case is when both estimates are highly uncertain (Figure 2d), as the risk of choosing the expected optimal plan increases further. This example highlights the importance of taking into account the distributions of the cost estimates rather than solely the expected values. $\square$
@@ 3.2.1. Uncertainty Modeling in ML
(3) $\begin{split}\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}=\mathcal{X}^{*})&=E[\text{Var}(f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\mathcal{D})]\\ &\quad+\text{Var}(E[f_{\theta}(\mathcal{X})|\mathcal{X}^{*},\mathcal{D}])\end{split}$
The first term in Equation 3 is aleatoric (or data) uncertainty and the second term is epistemic (or model) uncertainty (Kendall and Gal, 2017). The variability of the behavior of the real-world phenomena is explained by data uncertainty. Data uncertainty can also be caused by noisy input vectors or labels, as well as low-dimensional input vectors that do not adequately explain the sample (Hüllermeier and Waegeman, 2021). Model uncertainty captures the variability of predictions due to lack of knowledge about an ideal model. This is rooted in limitations in model architecture, training procedure, or the coverage of the training data (Gawlikowski, 2023). For a learned cost model, these notions can be quantified using various techniques from the ML literature (Gawlikowski, 2023). In the following, we explain our approach to quantify these two types of uncertainty.
Data and Model Uncertainties A neural network can be designed to predict the parameters of probability distribution (Nix and Weigend, 1994). i.e. it not only predicts the conditional expected value but also the conditional variance of the target given the training data and the input sample ($E[Var(y|x^{*},D)]$ with $y=f_{\theta}(\mathcal{X})$). This is done by adding a second branch to the output of the neural network, which predicts the variance. Optimizing the following loss function corresponds to maximizing the log-likelihood with a Gaussian prior (Nix and Weigend, 1994):
(4) $\text{Loss}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\ln(\sigma_{i}^{2})}{2}+\frac{(y_{i}-\mu_{i})^{2}}{2\sigma_{i}^{2}}+\frac{1}{2}\ln 2\pi\right)\quad$
where $N$ is the number of samples in $D$, $\mu_{i}$ is plugged in from the first branch of the output layer, $\sigma_{i}^{2}$ from the second one, and $y_{i}$ from the labels.
Model uncertainty is captured by the variance of the conditional expected values of the target ($Var(E[(y|x^{*},D)])$). To capture this, we use approximate variational inference by Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016). In this method, the neural network is trained by applying dropout on hidden layers with a Bernoulli distribution. At inference time, $T$ predictions are made with dropout applied randomly to the hidden layers, resulting in variations in the weight matrices, hence in the $T$ predicted values. With a Gaussian prior, this corresponds to combining multiple Gaussian distributions $\{N(\mu_{\hat{\theta}_{t}},\sigma_{\hat{\theta}_{t}}^{2})\}_{t=1}^{T}$ into one $N(\hat{\mu},\hat{\sigma}^{2})$. Given each draw of model weights results in two estimated parameters, we have a set of parameter estimates $\{(\hat{\mu}_{t},\hat{\sigma}_{t}^{2})\}_{t=1}^{T}$. The predicted mean is:
@@
$\hat{\sigma}_{m}^{2}=\text{Var}\left(E[(y|x^{*},D)]\right)=E\left[E\left[(y|x^{*},D)\right]^{2}\right]-\left(E\left[(y|x^{*},D)\right]\right)^{2}\$
(7) $=\frac{1}{T}\sum_{t=1}^{T}\hat{\mu}_{t}^{2}-\left(\frac{1}{T}\sum_{t=1}^{T}\hat{\mu}_{t}\right)^{2}$